OpenStudy (anonymous):

MATH HELP!!! 4|4-3X|=4x+6 how do i solve this?? will give medal!

OpenStudy (anonymous):

4(4-3x)=4x+6 16-12x=4x+6 10=16x x=1.6 just asking for sure, is the equation read like this : [4\times|4 - 3x|=4x+6\]

OpenStudy (anonymous):

@Leong there should be another answer like 1.6 or ________

OpenStudy (anonymous):

what are the answers for this question

OpenStudy (owlcoffee):

Whenever you see an absolute value that takes a fraction of the equation that includes the variable, in this case being \(\left| 4-3x \right|\) the "absolute value" part if the equation. "Absolute value" is defined as: \[\left| x \right| <=> \forall x \in /x \ge 0 ; \left| x \right| =x\] \[\left| x \right| <=> \forall x \in /x < 0 ; \left| x \right| =-x \] Now, what does this mean?, it means that when we deal with an absolute value we have to take in consideration two scenarios, the positive and the negative, so returning to the problem: \[4 \left| 4-3x \right|=4x+6\] We have to apply the definition of "absolute value" in order to solve, and to get rid if it, of course, so we will take the two scenarios: \(4 (4x-3)=4x+6\) and \((-4)(4x-3)=4x+6\). So you have to solve both of those equations in order to find the values of "x".

OpenStudy (anonymous):

oh yeah... sr, for got, then you just put a (-) => -4(4-3x)=4x+6 -16+12x=4x+6 8x=22 x=2.75